OpenStudy (anonymous):

Evaluate: 30 - m/6 for m = 30. Answer 25 0 6 1

OpenStudy (anonymous):

\[30- \frac{ m }{ 6 }\] \[m=30\] \[30-\frac{ 30 }{ 6 } = ?\]

OpenStudy (anonymous):

\[30 - 5 = 25\]